### v2

# Harmonizing Visual Representations for Unified Multimodal Understanding and Generation

Abstract
Unifying visual understanding and generation within a single multimodal framework remains a significant challenge, as the two inherently heterogeneous tasks require representations at different levels of granularity. Current approaches that utilize vector quantization (VQ) or variational autoencoders (VAE) for unified visual representation prioritize intrinsic imagery features over semantics, compromising understanding performance. In this work, we take inspiration from masked image modelling (MIM) that learns rich semantics via a mask-and-reconstruct pre-training and its successful extension to masked autoregressive (MAR) image generation. A preliminary study on the MAR encoder’s representation reveals exceptional linear probing accuracy and precise feature response to visual concepts, which indicates MAR’s potential for visual understanding tasks beyond its original generation role. Based on these insights, we present Harmon, a unified autoregressive framework that harmonizes understanding and generation tasks with a shared MAR encoder. Through a three-stage training procedure that progressively optimizes understanding and generation capabilities, Harmon achieves state-of-the-art image generation results on the GenEval, MJHQ30K and WISE benchmarks while matching the performance of methods with dedicated semantic encoders (e.g., Janus) on image understanding benchmarks. Our code and models will be available at https://github.com/wusize/Harmon.
Figure 1: An overview of image generation and understanding examples. All results are obtained by the proposed Harmon-1.5B, which uses a shared visual encoder for both tasks.
1 Introduction
What I cannot create, I do not understand.
— Richard Feynman
Generating images from textual descriptions and understanding images through targeted questions represent the forefront of AI research. With significant progress in both tasks, a plethora of text-to-image generation models [71, 9, 64, 100, 11, 69, 70, 3] and multimodal understanding LLMs [99, 53, 52, 44, 95, 21, 92, 17, 16, 88] have been developed, driven by advancements in model architectures and computation scaling. To push the envelope even further, unifying these two tasks into a single and cohesive framework that handles both understanding and generation is highly desirable for next-generation multimodal intelligence.
Early attempts [77, 78, 81, 84] have combined state-of-the-art diffusion models and multimodal LLMs by prompting the diffusion generators with LLM embeddings. These approaches are constrained by weak integration between image generation and text sequence modeling, resulting in suboptimal performance in instruction-based generation [32]. More recently, a tighter coupling of generation and understanding has been achieved through either next-token prediction approaches [54, 79, 89, 86] or denoising diffusion frameworks [98, 91, 97, 90], where visual representations for both tasks are jointly modeled.
Figure 2: Linear probing and feature visualization (using GradCAM++ [8]) provide crucial insights into the suitability of representations learned by VQGAN, VAE, and MAR models for serving as the visual encoder in a unified multimodal framework.
Visual understanding and generation are inherently heterogeneous tasks that require representations at different levels of granularity. In particular, visual understanding demands coarse-grained, high-level representations, whilst generation prefers fine-grained, intrinsic imagery features. Consequently, following standard practices in visual generation [71, 76], most approaches employ either a vector quantization (VQ) model or a variational autoencoder (VAE) to compress visual inputs into discrete tokens or continuous latents. These encoders are primarily pre-trained to preserve intrinsic image features for pixel-level reconstruction rather than capturing visual semantics, resulting in limited image understanding capabilities. Therefore, to obtain a harmonious representation with shared visual encoding for unified multimodal understanding and generation, a more integrated and moderate approach is warranted.
We take inspiration from masked image modeling (MIM) [34, 2] that can effectively encourage the model to develop a richer semantic representation of visual data via a simple mask-and-reconstruct pretext task. This learning by generating philosophy can learn balanced representations capable of both generation and understanding, making it a promising approach to reconcile these heterogeneous tasks in shared encoding. Moreover, the recent success of masked autoregressive (MAR) [46, 28] modeling that extends MIM to autoregressive image generation further substantiates the potential of MIM-based paradigms in this realm. It aligns with the de facto autoregressive prediction paradigm in MLLMs, making it more suitable for integrated synergy between generation and understanding.
Driven by this revelation, we conduct preliminary studies involving linear probing and feature visualization to measure the representation power of widely used VQGAN and VAE encoders in generative models and the MAR encoder, which is shown in Figure 2. Our findings reveal that VQGAN and VAE encoders struggle to capture high-level semantics, such as object concepts and attributes, which are crucial for effective visual understanding. In contrast, the MIM-trained MAR encoder achieves competitive linear probing accuracy and produces precise feature map activations. This verifies our hypothesis that a MIM-based approach can yield more robust semantic representations, which are vital for unifying the tasks of generation and understanding.
Figure 3: We propose Harmon, a unified framework with a shared MAR encoder for image generation and understanding tasks.
Building on these insights, in this paper, we present Harmon, a unified autoregressive model for multimodal generation and understanding as shown in Figure 3. To harmonize visual representations for two tasks in shared encoding, we propose a unified visual encoder (MAR-based) that captures both coarse semantics and fine-grained features and promotes consistent semantic grounding for both tasks. In Harmon, images are generated in a masked autoregressive manner while texts are regressed by next-token prediction. Furthermore, to align this visual encoder with the LLM language space, we design a three-stage training pipeline to gradually develop a unified multimodal generation and understanding model. Thanks to the versatility of the MIM paradigm, the proposed MAR-based encoder can be thoroughly optimized for both generation and understanding in each training stage.
We comprehensively assess Harmon on multimodal generation and understanding benchmarks. For text-to-image generation, it surpasses all unified methods with similar model scales on the GenEval benchmark [32], which measures the alignment between generated images and user instructions. Harmon also beats all unified models on the WISE benchmark [61] that necessitates world knowledge comprehension, Harmon produces images with higher visual quality, achieving state-of-the-art performance on the MJHQ30K benchmark [45]. For image understanding benchmarks, Harmon performs competitively with methods adopting a separate semantic encoder for visual understanding (i.e., Janus [87] and Janus-Pro [15]) while surpassing methods that use VQ or VAE encoders by a large margin. Moreover, the synergy between the two tasks via a shared representation is also observed in our experiment, where image generation is improved by co-training with image understanding.
2 Related Work
Multimodal Understanding. By injecting visual signals from a semantic encoder to an LLM, multimodal LLMs (MLLMs) [99, 21, 53, 52, 44, 57, 16, 85] demonstrate exceptional general and fine-grained image understanding capabilities. Such semantic encoders, including CLIP [68] and SigLIP [96], are usually pre-trained via contrastive vision-language aligning, a process where comprehensive visual concepts are learned. Despite their powerful visual understanding capabilities, these MLLMs are limited to image-conditioned question-answering without the ability to produce visual outputs.
Image Generation. Diffusion models [69, 71, 70, 64, 11, 3, 63, 51] have become the dominant paradigm for image generation, producing high-quality visual contents conditioned on class labels or language descriptions. Typically, the diffusion process is instantiated in the latent space of VAE [83, 40] for efficient training and inference. With the advancement of vision transformers, remarkable progress has also been made in autoregressive generation [26, 76, 93, 58] and masked generative models [5, 6, 1]. In these frameworks, images are usually compressed into discrete tokens using a VQ model, e.g., VQGAN [26]. More recently, the MAR [46] paradigm is introduced, following an encoder-decoder architecture and leveraging masked image modelling [34] for image generation.
Unified Visual Generation & Understanding. Pioneering works [77, 78, 81, 84] usually combine state-of-the-art diffusion models and multimodal LLMs, prompting the diffusion generators with LLM embeddings. Such paradigm is limited by the lack of deep interaction between image generation and text sequence modeling, exhibiting sub-optimal performance in instruction-based generation [32]. Besides, a few studies unify generation and understanding tasks as next token prediction [54, 79, 89, 86], where a VQGAN [26] compresses images into discrete tokens. There are also studies [98, 91, 97, 90] that regard LLMs as diffusion backbones, de-noising corrupted VAE latents for image generation but predicting next tokens for text regression. These two frameworks encode images using either VQGAN or VAE, which is a compression model by design, leading to diminution in visual understanding capability.
To address the above dilemma, ViLA-U [89] builds a unified vision tower by learning vector quantization on visual foundation models with joint contrastive text alignment and image reconstruction objectives. However, it struggles to balance the intricate interplay between semantic alignment and pixel-level fidelity. There are also studies that disentangle the forward pathways of understanding and generation. LlamaFusion [74] learns extra model weights that are only activated for vision signals. Likewise, Janus [87, 15] employs separate visual encoders for the two tasks, i.e., VQGAN [26] for generation and SigLIP [96] for understanding. Despite achieving state-of-the-art performances, this ad-hoc design neglects the potential of a unified representation that fosters deeper cross-modal synergy across both tasks. In contrast, we find that the MIM-trained MAR encoder naturally learns a representation well-suited for both generation and understanding. Based on this insight, we introduce Harmon, a unified framework that employs a shared encoder for both tasks.
Figure 4: The overall framework of Harmon. (a) Image generation is performed in the masked autoregressive manner. (b) Image understanding is formulated as image-conditioned text autoregression. The MAR encoder is shared by both tasks. Refer to caption represents the buffer embedding and Refer to caption represents the mask embedding.
3 Method
In this section, we introduce Harmon, a unified autoregressive framework comprising an MAR encoder, an LLM, and an MAR decoder as shown in Figure 4. For brevity, we omit the connection layers between the LLM and the MAR encoder/decoder.
3.1 Preliminaries
Masked Autoregressive Model. MAR [46] is a masked autoregressive model for image generation. Our preliminary results in Figure 2 suggest that the MAR encoder actually captures visual concepts and semantics during its generative pre-training, which are critical for visual understanding. This makes it a viable choice as an encoder to handle both visual generation and understanding within a single, unified framework beyond its original role.
In an MAR forward pass where $m$ patches are masked from an image ${\mathbf{X}}_{\text{img}}$ with height $h$ and width $w$, the encoder $f_{\text{enc}}$ takes the $hw-m$ seen image patches ${\mathbf{X}}_{\text{seen}}$ and $n$ buffer embeddings ${\mathbf{X}}_{\text{buffer}}$ as input for feature extraction
${\mathbf{Z}}_{\text{enc}}=f_{\text{enc}}({\mathbf{X}}_{\text{seen}},{\mathbf{X}}_{\text{buffer}}).$ (1)
Then ${\mathbf{Z}}_{\text{enc}}$ outputted by the encoder and a set of mask embeddings ${\mathbf{Z}}_{\text{mask}}$ are fed into the decoder $f_{\text{enc}}$, which predicts the $m$ masked image patches
${\mathbf{X}}_{\text{mask}}=f_{\text{dec}}({\mathbf{Z}}_{\text{mask}},{\mathbf{Z}}_{\text{enc}}).$ (2)
Large Language Model. A modern large language model (LLM) typically follows GPT [66, 67, 4] to adopt a decoder-only architecture with causal attention. In a forward pass, a sequence of text tokens are first mapped into text embeddings and then fed into the LLM $f_{\text{LLM}}$. Each token embedding interacts with preceding ones in the LLM’s causal attentions and predicts the next token by a text head on top of the LLM.
3.2 Text-to-Image Generation
Image generation in Harmon works in a masked autoregressive manner as illustrated by Figure 4(a).
[…]
Encoder to LLM. As shown in Eq. 1 and Figure 4(a), the output of the MAR encoder ${\mathbf{Z}}_{\text{enc}}$ includes image embeddings of seen patches and buffer embeddings. Instead of directly feeding ${\mathbf{Z}}_{\text{enc}}$ to the MAR decoder for reconstruction, we pass ${\mathbf{Z}}_{\text{enc}}$ to the LLM following the prompt text embeddings ${\mathbf{Z}}_{\text{txt}}$ for multimodal integration.
LLM to Decoder. The output produced by the LLM consists of the processed prompt embeddings and MAR encoder outputs. We extract the updated encoder outputs ${\mathbf{Z}}^{\prime}_{\text{enc}}$, and feed them to the MAR decoder $f_{\text{dec}}$ together with the $m$ mask embeddings ${\mathbf{Z}}_{\text{mask}}$. Then, $f_{\text{dec}}$ predicts the masked image patches
${\mathbf{X}}_{\text{mask}}=f_{\text{dec}}({\mathbf{Z}}_{\text{mask}},{\mathbf{Z}}^{\prime}_{\text{enc}}).$ (3)
Since ${\mathbf{Z}}^{\prime}_{\text{enc}}$ has interacted with text embeddings in the LLM, the prediction is conditioned on the language prompts.
Table 1: Detailed hyperparameters in training. Data ratio refers to the ratio of text data, image-to-text data, and text-to-image data.
Setting Stage I Stage II Stage III
MAR Decoder: $10^{-4}$ MAR Decoder: $10^{-4}$ MAR Decoder: $2\times 10^{-5}$
LR. MAR Encoder: $10^{-5}$ LLM & MAR Encoder: $10^{-5}$ LLM & MAR Encoder: $2\times 10^{-6}$
Image Size 256 256 512
Batch Size 4096 4096 1024
Data Ratio 0:1:2 1:3:8 1:3:16
Training Step 50,000 50,000 50,000
Training Objective. We follow MAR [46] to reparameterize the generation of masked patches as a reverse diffusion process [36]. Therefore, the training loss for predicting a masked patch $x_{\text{mask}}\in{\mathbf{X}}_{\text{mask}}$ is
$\mathcal{L}(x_{\text{mask}},x_{\text{gt}})=\mathbb{E}_{\varepsilon,t}\left[\left\|\varepsilon-\varepsilon_{\theta}(x_{t}|t,x_{\text{mask}})\right\|^{2}\right].$ (4)
Here, $\varepsilon$ is a noise vector sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. $x_{t}$ is a noise-corrupted vector from $x_{\text{gt}}$ and $t$ is a randomly sampled time step of the noise schedule. The noise estimator $\varepsilon_{\theta}$ is a small MLP network defined in  [46].
Inference. Our masked autoregressive generation consists of $K$ forward passes, starting with the masked patch number $m_{0}=hw$, i.e., all image patches are masked and only the buffer embeddings are fed into the MAR encoder. We gradually reduce the mask ratio following a cosine curve: $m_{k}=hw\cdot\text{cos}(\frac{k}{2K}\pi)$. Here, $k$ means the $k$-th generation step. And the number of predicted image patches is $m_{k}-m_{k-1}$ in the $k$-th step. Since the attention in the LLM is causal, prompt embeddings are only passed to the LLM in the first iteration, leaving key-value caches for later steps. In addition, we also follow prior works [46, 91, 86] to employ the classifier-free guidance (CFG) [35].
3.3 Image Understanding
As shown in Figure 4(b), multimodal image understanding is formulated as question-answering conditioned on image contents, where the text autoregression is performed as next-token prediction.
[…]
Visual Encoding. Different from the text-to-image generation process, we pass all the image patches without any masking together with the buffer embeddings to the MAR encoder $f_{\text{enc}}$
${\mathbf{Z}}_{\text{enc}}=f_{\text{enc}}({\mathbf{X}}_{\text{img}},{\mathbf{X}}_{\text{buffer}}).$ (5)
Answer Generation. The encoder output ${\mathbf{Z}}_{\text{enc}}$ together with text embeddings in the prompt are fed to the LLM. The corresponding textual answer is generated by the LLM in an autoregressive manner, conditioned on the image and the question. A cross-entropy loss is applied to supervise the prediction of answer tokens during training.
3.4 Training Recipe
The training stages and the associated data sources play a pivotal role in unleashing the potential of the shared encoder in Harmon. We split the training of Harmon into three stages, progressively enhancing the model’s generation and understanding abilities. More information about the data sources is provided in the appendix.
Stage I: Vision-Language Alignment. Since the MAR encoder/decoder and the LLM undergo unimodal pre-training on images and texts, respectively, it is necessary to align these two modalities on a large corpus of image-text pairs. Specifically, we use 22M images with dense captions from public datasets [44, 75, 47, 49] for knowledge-rich training of the MAR encoder, enabling language-based image understanding in Harmon. To maintain the original generation capability of MAR, we employ ImageNet1K [23] with 1.2M data samples for class-conditional generation, treating class names as image captions. In this stage, only the MAR encoder and decoder are trainable.
Stage II: Comprehensive Multimodal Training. After aligning the MAR encoder/decoder with the LLM, we utilize diverse question-answering data and text-to-image data to enhance generation and understanding abilities, with the LLM unlocked. Specifically, we draw 20M samples from Infinity-MM [33] plus 5M dense caption samples from stage I for multimodal understanding. For text-to-image generation, we collect 50M images from public datasets [60, 59, 22, 75, 44, 47]. All of these images come with dense captions generated by LMMs [80, 12, 85], which are re-written by an LLM [82] for more concise prompts focusing on generation.
Stage III: High-Quality Fine-tuning. We adopt higher-quality data to fine-tune Harmon in stage III. For multimodal question-answering, we employ the smaller but more balanced instruction-tuning data from LLaVA-One-Vision [44]. For text-to-image generation, we filter the 50M images in stage II based on image sizes and aesthetic scores, with 20% of them reserved. In addition, we include 6M synthetic images [62, 38] to enhance visual quality. In stage III, we also increase the image resolution from 256 to 512 to strengthen both generation and understanding capabilities.
Table 2: Evaluation of image understanding capabilities on multimodal question-answering benchmarks. Und. Only stands for models trained for image understanding only. means employing a separate semantic encoder.
[…]
Und. Only LLaVA-v1.51.51.51.5-Phi-$1.5$ [91] CLIP ViT-L [68] 1.3B 84.1 1128 - - - 56.5 30.7
MobileVLM [19] CLIP ViT-L [68] 1.4B 84.5 1196 - 53.2 - 56.1 -
MobileVLM-V2 [20] CLIP ViT-L [68] 1.4B 84.3 1303 - 57.7 - 59.3 -
DeepSeekVL [57] SigLIP-Large [96] 1.3B 88.3 1307 225 64.6 - 59.3 33.8
MiniGemini [49] CLIP ViT-L [68] 2B 83.9 1341 312 59.8 - 59.9 -
&ConvNext-L [56]
Unified ILLUME [84] OpenCLIP ViT-H [18] 7B 88.5 1445 - 65.1 72.9 - 38.2
TokenFlow-XL [65] CLIP ViT-B(VQ) [65] 13B 86.8 1546 - 68.9 68.7 62.7 38.7
LWM [54] VQGAN [26] 7B 75.2 - - - - 44.8 -
VILA-U [89] SigLIP-Large(VQ) [89] 7B 85.8 1402 - 59.0 60.8 -
Chameleon [79] VQGAN [26] 7B - - - - - - 22.4
D-Dit [50] VAE [27] 2.0B 84.0 1125 - - - 59.2 -
Show-o [91] MAGVIT-v2 [93] 1.3B 80.0 1097 248 51.6 54.4 58.0 26.7
Janus [87] SigLIP-Large [96] 1.3B 87.0 1338 222 69.4 63.7 59.1 30.5
Janus-Pro [15] SigLIP-Large [96] 1.5B 86.2 1444 268 75.5 68.3 59.3 36.3
Harmon-0.5B MAR-B [46] 0.5B 86.5 1148 260 59.8 62.5 56.3 34.2
Harmon-1.5B MAR-H [46] 1.5B 87.6 1155 321 65.5 67.1 58.9 38.9
Table 3: Text-to-image generation on MSCOCO-30K and MJHQ-30K. FID is used as the metric for both benchmarks. Gen. Only stands for models trained for image generation only.
[…]
Gen. Only DALL·E 2 [70] $10.39$
GigaGAN [39] $9.09$
SD1.5 [72] $9.62$
PixArt-$\alpha$ [10] 7.32 9.85
SDXL [64] 7.38 8.76
SD2.1 [71] - 26.96
LlamaGen [76] - 25.59
Unified Show-o [91] 9.24 $15.18$
LWM [54] 12.68 $17.77$
VILA-U [89] - $7.69$
Janus [87] 8.53 $10.10$
Janus-Pro-1.5B [15] 16.08 9.53
Harmon-0.5B 8.86 6.08
Harmon-1.5B 8.39 5.15
4 Experiments
We conduct our experiments on two model variants, Harmon-0.5B and Harmon-1.5B. Harmon-0.5B is built on Qwen2.5-0.5B-Instruct [92] with visual encoder and decoder from MAR-B [46] while Harmon-1.5B is based on Qwen2.5-1.5B-Instruct [92] and MAR-H [46]. Image size is fixed at $256\times 256$ in stage I and stage II, which is increased to $512\times 512$ in stage III. All hyperparameters in the three-stage training are detailed in Table 1. The whole training process costs $4$ days and $8$ days for Harmon-0.5B and Harmon-1.5B, respectively, with $32$ Nvidia A100100100100 ($80$GB) GPUs. In inference, greedy search is employed for text generation in image understanding. For text-to-image generation, we use $K=64$ MAR forward passes and set CFG weight as 3.0.
4.1 Comparisons with Existing Models
Image Understanding. We evaluate our model’s image understanding capabilities on widely used multimodal question-answering benchmarks, including POPE [48], MME [29], MMB [55], SEED [43], GQA [37] and MMMU [94]. As shown in Table 2, we compare Harmon with both unified models and understanding-only models (indicated as Und. Only). Given the model size of Harmon, we mainly focus on multimodal models with LLM under 2B parameters. It is noteworthy that Harmon achieves performance on par with understanding-only models and Janus models, which employ a separate semantic encoder (i.e., SigLIP [96]) for visual perception. Besides, Harmon outperforms prior unified models based on VAE/VQGAN encoders, such as LWM [54], Chameleon [79], D-DiT [50] and Show-O [91].
Text-to-Image Generation. We assess both the visual quality and controllability of image generation. Specifically, we report Fréchet Inception Distance (FID) scores on the MSCOCO-$30$K [14] and MJHQ-$30$K [45] benchmarks, measuring the similarity between the distributions of generated images and reference images. Besides, we evaluate our models on the GenEval [32] benchmark, which examines how well the attributes of generated objects can be controlled by user instructions, including counting, position and colour attributes. As shown in Table 3, Harmon outperforms all unified models on MJHQ-$30$K that gauges the aesthetic quality of generated images. For evaluation of controllability on GenEval, Harmon outperforms all unified and generation-only models as shown in Table 4. Notably, our smaller variant Harmon-0.5B is already comparable to Janus-Pro-1.5B, indicating the superior vision-language alignment brought by the shared visual encoder. Additionally, we assess Harmon’s ability to understand complex semantic and world knowledge using the WISE benchmark [61], where implicit prompts like “Einstein’s favorite musical instrument” are provided. As shown in Table 5, Harmon archives the best performance among all unified models.
Table 4: Evaluation of text-to-image generation on GenEval benchmark. Gen. Only stands for models trained for image generation only.
[…]
Gen. Only LlamaGen [76] 0.71 0.34 0.21 0.58 0.07 0.04 0.32
LDM [72] 0.92 0.29 0.23 0.70 0.02 0.05 0.37
SDv1.5 [72] 0.97 0.38 0.35 0.76 0.04 0.06 0.43
PixArt-$\alpha$ [10] 0.98 0.50 0.44 0.80 0.08 0.07 0.48
SDv2.1 [72] 0.98 0.51 0.44 0.85 0.07 0.17 0.50
DALL-E 2 [70] 0.94 0.66 0.49 0.77 0.10 0.19 0.52
Emu3-Gen  [86] 0.98 0.71 0.34 0.81 0.17 0.21 0.54
SDXL [64] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 [3] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
SD3-Medium [27] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
Unified LWM [54] 0.93 0.41 0.46 0.79 0.09 0.15 0.47
SEED-X [31] 0.97 0.58 0.26 0.80 0.19 0.14 0.49
Show-o [91] 0.95 0.52 0.49 0.82 0.11 0.28 0.53
D-DiT [50] 0.97 0.80 0.54 0.76 0.32 0.50 0.65
Transfusion [98] - - - - - - 0.63
ILLUME [84] 0.99 0.86 0.45 0.71 0.39 0.28 0.61
TokenFlow-XL [54] 0.95 0.60 0.41 0.81 0.16 0.24 0.55
OmniGen [90] 0.99 0.86 0.64 0.85 0.31 0.55 0.70
Chameleon [79] - - - - - - 0.39
Janus [87] 0.97 0.68 0.30 0.84 0.46 0.42 0.61
Janus-Pro-1.5B [15] 0.98 0.82 0.51 0.89 0.65 0.56 0.73
Harmon-0.5B 0.99 0.80 0.57 0.87 0.55 0.48 0.71
Harmon-1.5B 0.99 0.86 0.66 0.85 0.74 0.48 0.76
Table 5: Evaluation of text-to-image generation on WISE benchmark. Gen. Only stands for models trained for image generation only.
[…]
Gen. Only SDv1.5 [72] 0.34 0.35 0.32 0.28 0.29 0.21 0.32
SDv2.1 [72] 0.30 0.38 0.35 0.33 0.34 0.21 0.32
Emu3-Gen  [86] 0.34 0.45 0.48 0.41 0.45 0.27 0.39
FLUX.1-schnell [42] 0.39 0.44 0.50 0.31 0.44 0.26 0.40
SD3-Medium [27] 0.42 0.44 0.48 0.39 0.47 0.29 0.42
SDXL [64] 0.43 0.48 0.47 0.44 0.45 0.27 0.43
SD3.5-Large [27] 0.44 0.50 0.58 0.44 0.52 0.31 0.46
PixArt-$\alpha$ [10] 0.45 0.50 0.48 0.49 0.56 0.34 0.47
FLUX.1-dev [42] 0.48 0.58 0.62 0.42 0.51 0.35 0.50
Unified Janus [87] 0.16 0.26 0.35 0.28 0.30 0.14 0.23
Janus-Pro-1.5B [15] 0.20 0.28 0.45 0.24 0.32 0.16 0.26
Orthus [41] 0.23 0.31 0.38 0.28 0.31 0.20 0.27
VILA-U [89] 0.26 0.33 0.37 0.35 0.39 0.23 0.31
Show-o [91] 0.28 0.40 0.48 0.30 0.46 0.30 0.35
Harmon-1.5B 0.38 0.48 0.52 0.37 0.44 0.29 0.41
Qualitative Comparison. In Figure 5, we visually compare our generation results with state-of-the-art unified models, i.e., Show-o [91] and Janus-Pro [15]. For a fair comparison, we select Show-o-1.3B-512, Janus-Pro-1.5B and our Harmon-1.5B. We observe better consistency between images and prompts in the results of Harmon. For example, Harmon is able to precisely generate the ‘pink stop sign’ while Show-o and Janus-Pro fail to generate the textual contents. Harmon also successfully handles the ‘sink’ that is missed in Show-o’s generation result. Moreover, Harmon tends to produce images of higher visual quality, such as the ‘paper artwork’ and ‘dog’ in the bottom left and bottom middle examples, where Show-o and Janus-Pro generate distorted results. We provide more generation results and comparisons with more models in the appendix.
Figure 5: Qualitative comparison between Show-o (1.3B), Janus-Pro (1.5B) and our Harmon-1.5B on text-to-image generation. The text below each image represents the generation prompt, with key terms guiding the generation highlighted in orange. Best viewed on screen.
4.2 Analysis & Ablation
In this section, we study the effectiveness of our design choices. All experiments are based on Harmon-0.5B.
Encoders for Visual Understanding. We examine the potential of visual encoders for understanding tasks, including VQGAN, VAE and MAR, which are trained by generative losses. Additonally, we include a semantic encoder, SigLIP [96], which is trained by contrastive vision-language aligning. For a fair comparison, MAR-B and SigLIP-Base-Patch16 are chosen. As shown in Table 8, VQGAN and VAE, which focus on image intrinsic features like local details and textures for high-fidelity compression, lag behind MAR and SigLIP by large margins. Meanwhile, MAR encoder performs comparably with SigLip, a language-aligned semantic encoder. The results are consistent with findings in MAE [34] that visual encoders can learn essential visual semantics and concepts through mask-ratio MiM training. In addition, the benchmark results also conform with the linear probing accuracies observed in our preliminary study (Figure 2).
Table 6: Choices of visual encoders for understanding. ‘Acc’ represents the top-5 linear probing accuracy on ImageNet [23].
[…]
1 SigLIP 95.9 85.1 1203 258 61.1 63.2 56.1 35.2
2 VQGAN 18.2 57.2 67.3 21.8 37.3 38.3 38.0 27.7
3 VAE 18.5 63.8 732 223 44.9 42.5 40.2 30.3
4 MAR 83.1 86.1 1123 262 60.1 62.2 55.7 33.3
Table 7: Ablation study on the effects of the three training stages.
[…]
1 2 3
1 85.2 1003 218 51.4 33.2 10.23 7.56 0.66
2 84.7 1064 217 50.2 32.8 10.99 8.12 0.65
3 85.6 $1$ 251 54.1 34.0 15.64 16.85 0.56
4 86.5 1148 260 56.3 34.2 8.86 6.08 0.71
Table 8: Ablation study on the effect of image resolution in stage III.
[…]
1 256 86.1 1120 258 55.4 32.6 11.50 9.91 0.68
2 384 86.5 1144 260 55.5 33.7 10.97 8.65 0.69
3 512 86.5 1148 260 56.3 34.2 8.86 6.08 0.71
Training Stages. In Table 8, we study the impact of each training stage. By comparing rows #1, #2 & #3 with row #4, we observe that all training stages contribute to the overall performance. Specifically, the alignment in stage I (#1 and #4) and comprehensive training in stage II (#2 and #4) mainly affect the performance on understanding benchmarks while the high-quality training in stage III, which features aesthetical images with higher resolutions (#3 and #4), improves generation performances the most, especially on the MJHQ-30K benchmark that measures visual quality.
Image Resolutions. In Table 8, we study the effect of increasing image resolution in the final training stage. Overall, performances on both understanding and generation tasks increase with higher resolutions. Specifically, the MSCOCO-30K and MJHQ-30K benchmarks that demand visual quality are more sensitive to image resolutions, where FID values drop drastically when image resolution increases. In contrast, performance gains are less pronounced on the general image understanding tasks as well as the GenEval benchmark that weighs more on alignment between images and texts in instructed-based generations.
Figure 6: Ablation study on the mutual effect between understanding and generation supervision. Our shared encoding design brings performance gains on image generation (b) while maintaining understanding capability (a).
Single Task v.s. Unified Training. We study the mutual effect between generation and understanding by training with single-task losses. As shown in Figure 6(a), the unified training of the two tasks does not decrease scores on understanding tasks compared with single-task training. For image generation, we observe that the understanding loss even boosts generation performance, especially on the GenEval benchmark that measures the consistency between generated images and user instructions, as suggested by Figure 6(b). In addition, we also compare our shared encoding strategy with a Janus-like design, i.e., using a semantic encoder (SigLIP-Base-Patch16 [96]) for understanding and MAR encoder for generation. As shown in Figure 6 (c,d), the decoupled encoding architecture accommodates the two tasks well. However, the synergistic effect disappears, suggesting that a shared representation is key to the co-evolution of understanding and generation capabilities within a unified model.
5 Conclusion
In this paper, we investigate the visual encoders in unified understanding and generation frameworks. Through a preliminary study including linear probing and feature visualization, we recognize the potential of MAR for visual understanding beyond its in-born generative role. Based on this finding, we present Harmon, a unified framework that harmonizes image understanding and generation in MAR’s encoding space. Through a progressive three-stage training procedure, Harmon obtains state-of-the-art results on text-to-image generation benchmarks while achieving competitive performance on multimodal question-answering benchmarks. Moreover, we observe the synergy between the two tasks in our experiment, where the co-training with understanding loss boosts generation performance.
Acknowledgement.
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-PhD-2023-08-048T), the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).